# A Unified Framework for Direction and Diffuseness Estimation Using Tight-Frame Microphone Arrays

### v2

@@ III.3 Evaluation metrics
Intensity–Energy ratio $\Psi_{\mathrm{IE}},\>\Psi_{\mathrm{AVE}}$
From pressure $p$ and velocity $\bm{u}$ (via A-format or HOA/FOA processing) in the cases of Afmt and Fibo64, the diffuseness $\Psi_{\mathrm{IE}}$ is defined as
@@
For TF24 case, we compute directional $\Psi_{r_{i}}$ defined as Eq. (19). In directions where only uncorrelated background noise is observed, the estimated diffuseness tends to assume relatively high values. However, such regions do not correspond to acoustically significant signals. This observation suggests the necessity of introducing directional weighting.
One possible choice is to weight the diffuseness by the energy associated with each direction. After several preliminary trials, weighting by the squared magnitude of the particle velocity was adopted, as it exhibited higher sensitivity to spatial variations in the sound field. Fig. 7 illustrates an example of the directional energy distribution and the corresponding weighting factors based on the squared particle velocity, computed for Case 2. The incident beam is assumed to arrive predominantly from the 9th direction, and it can be observed that the velocity-based weighting responds subtly but consistently to this directional bias.
@@
Figure 7: Example of weighting applied to the directional diffuseness: results for Case 2 with a beam ratio of $\eta=0.95$ at 2 kHz.
@@
In the present study, however, we emphasize a velocity-only route and form
@@
which directly captures the spatial anisotropy of the local particle-velocity distribution without explicit reference to the pressure term.
This choice is first motivated by the characteristics of the Tight frame array. Since each microphone in the TF24 configuration is a highly directional (DPA4017 in our case) rather than an ideal cardioid, the sum of each microphone pair does not accurately reproduce the omnidirectional pressure component.
Consequently, defining diffuseness in terms of pressure-velocity coherence becomes less reliable in TF24. By constructing the covariance matrix solely from the velocity vectors, we can evaluate the spatial uniformity in a physically consistent and array-independent manner, thus ensuring the robustness of diffuseness estimation even under non-ideal directivity conditions.
@@
In the TF24 system, the particle velocity components used to construct the covariance matrix $\bm{C}_{uu}$ are obtained by projecting the velocity vectors $u_{\bm{r}_{i}}$, measured in 12 directions, onto the global x,yx,\;y, and $z$ axes. This projection is performed by multiplying each directional component by the corresponding direction vector $\bm{R}$ as,
$\bm{u}=\begin{bmatrix}u_{x}\\ u_{y}\\ u_{z}\end{bmatrix}=\frac{1}{4}\bm{R}^{\!\top}\bm{u}_{\bm{r}},$ (26)
where $\bm{u}_{\bm{r}}=\left[u_{\bm{r}_{1}}u_{\bm{r}_{2}}\cdots u_{\bm{r}_{12}}\right]^{\!\top}$.
@@ IV Results and discussion
Figure 8: Case 1: Diffuseness indices based on the I/E method: for Afmt and Fibo64, the results show $\Psi_{\mathrm{IE}}$; for TF24, $\Psi_{\mathrm{AVE}}$ is presented.
The Afmt results exhibited relatively large values in the low-frequency region and above 4 kHz, while the TF24 array also showed slightly elevated values of approximately 0.2 at 16 kHz. In contrast, the Fibo64 array, which employs 64 microphones, yielded values close to zero across all frequency bands.
For the eigenvalue-based indices, although the absolute magnitudes differed across methods, all 2,520 incident directions yielded values of zero up to three decimal places when averaged over the nine octave bands from 63 Hz to 16 kHz.
@@
Directional accuracy of the active-intensity vector was also compared. The results are shown in Fig. 9. Although diffuseness and the angular error in intensity detection are entirely different quantities, the overall trend of the results is similar to that shown in Fig. 8. For the Afmt system, the estimation accuracy is low in both the low- and high-frequency ranges, and the frequency band in which the correct incident direction can be detected is limited to approximately 500 Hz to 2 kHz. In contrast, the Fibo64 system achieves almost error-free direction detection across the entire frequency range. The performance of TF24 lies between these two cases, showing an average detection error of about 4° at 8 kHz and slightly less than 6° at 16 kHz. Considering that these values represent the average error over 2520 incident directions, the overall performance can be regarded as practically sufficient.
Figure 9: Case 1: Mean detection error of the incident angle estimated from acoustic intensity obtained by each microphone array: Average detection error over 2520 incident directions.
@@
Figure 10 shows the results of the COMEDIE-based diffuseness index, $\Psi_{\mathrm{COM}}$, for the TF24 system. Although slight fluctuations can be observed in the high-frequency range, the results closely follow the theoretical value of $1-\eta$ from the low to mid frequencies. All indices exhibit an approximately linear relationship, and the correlation between $1-\eta$ and the diffuseness estimated in each frequency band is very high. Because the indices are therefore difficult to distinguish in direct comparison, we quantified their deviation more explicitly by calculating the absolute difference between the diffuseness value and $1-\eta$, and taking the maximum of this difference for each index. The results are shown in Fig. 11.
In Fig. 11, the results obtained using the intensity/energy (I/E)-based indices are shown as solid lines, while those based on the eigenvalue methods—particularly COMEDIE—are shown as dashed lines. The eigenvalue-based results show only small deviations across all frequency bands. A similar tendency to that seen in Figs. 8 and 9 is observed for the Afmt system, especially in the I/E-based indices, indicating that the frequency range with small estimation error is relatively narrow. In contrast, even for Afmt, the indices derived from the eigenvalues of the velocity covariance matrix exhibit comparatively small errors.
Figure 10: Case 2: Example of the variation in diffuseness when the beam ratio $\eta$ is varied from 0 to 1 in steps of 0.05. Results are obtained for the TF24 array using the eigenvalues of the covariance matrix and the COMEDIE, i.e., $\Psi_{\mathrm{COM}}$. The numbers in the figure indicate $\eta$; in the ideal case, the diffuseness corresponds to $1-\eta$.
Figure 11: Case 2: Maximum absolute error between the beam ratio $1-\eta$ and the diffuseness, calculated with I/E (solid lines) and eigenvalue examination; $\Psi_{\mathrm{COM}}$ (broken lines).
@@
Figure 12: Case 3: Diffuseness for the interference field formed by two plane waves at various incident angles at $1\text{\,}\mathrm{k}\mathrm{H}\mathrm{z}$. (a) Afmt, (b) Fibo64, and (c) TF24. The results for $\Psi_{\mathrm{IE}}$, $\Psi_{\mathrm{PR}}$, and $\Psi_{\mathrm{COM}}$ are shown. For reference, in (a), the results of $\Psi_{\mathrm{PR}}$ and $\Psi_{\mathrm{COM}}$ obtained using all four FOA components ($p$ and $\bm{u}$) are also presented. In (b), the results of $\Psi_{\mathrm{COM}}$ calculated from the HOA coefficients (up to 4th) are additionally shown.
@@
Figure 13: Diffuseness calculated from waveforms obtained by mixing impulse responses measured in anechoic and reverberation rooms at various energy ratios. The energy ratio of the anechoic component was varied from 0 to 1 in 0.1 increments. The left panels show the results for the Afmt array, and the right panels show those for the TF24 array. The upper panels correspond to the intensity-energy (I/E) method, while the lower panels show the eigenvalue-based results using the COMEDIE index.
@@ V.1 Concept
Finally, the trends obtained from the simulations were experimentally verified using measured data. Two sound fields were considered: an anechoic room and a reverberation room. As expected, diffuseness is nearly zero in the anechoic room and close to unity in the reverberation room. However, unlike the numerical simulations, neither environment can be regarded as perfectly ideal, and it is difficult to ensure that the measured results precisely match these theoretical values. Therefore, to enable a meaningful verification, we adopted an approach analogous to Case 2 of the benchmark simulations: the impulse responses measured in the anechoic and reverberation rooms were used to represent the “beam” and “diffuse” components, respectively. By mixing these two responses at arbitrary energy ratios, we examined the linearity between the prescribed ratio and the resulting diffuseness.
@@ V.2 Conditions
For the Afmt array, an AMBEO VR Mic was used, and for the TF24 array, DPA 4017 microphones were employed with a spacing of 20 mm between each microphone pair. As a preliminary measurement conducted in an anechoic room to determine the acoustic reference point of the shotgun microphone, impulse responses were measured twice at the same position: once with the microphone oriented toward the sound source, and once with it oriented perpendicularly. The time difference between the peaks of these two responses was used to estimate the reference point, which was found to be approximately 61 mm from the bottom of the microphone (amplifier section). Using this reference, the array was mounted on a rotatable jig (a camera tripod head) that allowed precise rotations in 45° increments around a point located 10 mm above the estimated center.
@@ V.3 Results
The results are shown in Fig. 13. The left panels correspond to the Afmt array, and the right panels to the TF24 array. The upper panels show the diffuseness obtained using the intensity-energy (I/E) method, while the lower panels present the diffuseness $\Psi_{\mathrm{COM}}$ calculated from the eigenvalues of the covariance matrix constructed only from the velocity components of the FOA signals. It is evident that the results for the Afmt array exhibit a larger deviation from $1-\eta$ in both low and high frequency ranges, and the overall trend appears distorted. In contrast, the TF24 array demonstrates a more linear relationship. The fact that the diffuseness does not reach exactly unity when $\eta=0$ (reverberation room only) and does not converge to zero when $\eta=1$ (anechoic room only) is attributed to both the non-ideal characteristics of the measured sound fields and the inevitable experimental errors, so the absolute values should be interpreted with appropriate caution.
Figure 14: The maximum absolute difference between the calculated diffuseness and $1-\eta$ for each index

### v3

@@ III.3 Evaluation metrics
Intensity-Energy ratio $\Psi_{\mathrm{IE}},\>\Psi_{\mathrm{AVE}}$, Coefficient-of-variation $\Psi_{\mathrm{CV}}$
From (pseudo) pressure and (pseudo) velocity (via A-format or HOA/FOA processing) in the cases of Afmt and Fibo64, the (pseudo) intensity $\bm{I}(f)$ and the (pseudo) energy density $E(f)$ are computed and the diffuseness $\Psi_{\mathrm{IE}}$ is defined as
@@
For TF24 case, we compute directional $\Psi_{r_{i}}$ defined as Eq. (19). In directions where the incident sound energy is negligible, the pseudo pressure and particle-velocity components estimated by the TF24 array become very small and are comparable to background or numerical noise. Because the diffuseness measure is defined as a normalized ratio, the residual uncorrelated components can become relatively dominant in such low-energy directions, leading to artificially high diffuseness values. Importantly, these regions do not correspond to acoustically meaningful sound-field contributions, but rather to directions with insufficient signal energy. This behavior indicates that directional weighting based on signal strength is necessary to suppress physically irrelevant directions when evaluating diffuseness.
After several preliminary trials, weighting by the squared magnitude of the particle velocity was adopted, as it exhibited higher sensitivity to spatial variations in the sound field. Fig. 7 illustrates an example of the directional energy distribution and the corresponding weighting factors based on the squared particle velocity, computed for Case 2. The incident beam is assumed to arrive predominantly from the 9th direction (9th row of the direction vector matrix $\bm{R}$ in Eq. (13), i.e., $x$-axis), and it can be observed that the velocity-based weighting responds subtly but consistently to this directional bias.
@@
Figure 7: Numerical simulation results of directional weighting applied to the diffuseness estimation. Results are shown for Case 2 with a beam ratio of $\eta=0.95$ at 2 kHz. In this condition, 95% of the acoustic energy is carried by a single plane wave arriving approximately from the $x$-axis direction. This dominant incidence corresponds to the 9th row of the direction vector matrix $\bm{R}$ in Eq. (13), and therefore the squared particle-velocity (or energy) associated with the 9th direction becomes significantly larger than those of the other directions.
Finally, in addition to the intensity-energy ratio based measures, we also consider a coefficient-of-variation-based diffuseness indicator, denoted as $\Psi_{\mathrm{CV}}$.
The coefficient-of-variation approach was originally introduced by Ahonen[16] and later employed in the context of spatial audio by Del Galdo et al.[17], where diffuseness is characterized by the statistical variability of acoustic intensity. In this study, $\Psi_{\mathrm{CV}}$ is formulated exclusively in terms of the acoustic intensity vector, without explicit use of the energy density.
For a set of realizations (or time–frequency samples), $\Psi_{\mathrm{CV}}$ is defined as
$\Psi_{\mathrm{CV}}=\sqrt{1-\frac{\|\mathbb{E}[\bm{I}(f)]\|}{\mathbb{E}[\|\bm{I}(f)\|]}},$ (25)
where $\bm{I}$ denotes the instantaneous acoustic intensity vector and $\mathbb{E}[\cdot]$ represents expectation over realizations. This formulation yields $\Psi_{\mathrm{CV}}\approx 0$ for a highly directional (traveling-wave) field and $\Psi_{\mathrm{CV}}\approx 1$ for an isotropic diffuse field.
Unlike $\Psi_{\mathrm{IE}}$ and $\Psi_{\mathrm{AVE}}$, which rely on the ratio between mean intensity and energy density, $\Psi_{\mathrm{CV}}$ quantifies diffuseness purely through the statistical dispersion of the intensity vector. In the present work, $\Psi_{\mathrm{CV}}$ is included as a reference intensity-based diffuseness measure, enabling direct comparison with the proposed directionally weighted diffuseness metrics and their behavior is specifically examined using benchmark Case 2 mentioned above.
@@
In the present study, however, we emphasize a velocity (including pseudo) only route and form
@@
which directly captures the spatial anisotropy of the local (pseudo)velocity distribution without an explicit reference to the pressure term.
This choice is first motivated by the characteristics of the Tight frame array. Since each microphone in the TF24 configuration uses a model of practical directional characteristics rather than an ideal cardioid, the sum of the outputs of each microphone pair does not necessarily reproduce the ideal omnidirectional pressure component.
Consequently, defining diffuseness in terms of pressure-velocity coherence may become less reliable in TF24. Instead, by constructing the covariance matrix solely from the velocity vectors, we can evaluate the spatial uniformity in a physically consistent and array-independent manner, thus ensuring the robustness of diffuseness estimation even under non-ideal directivity conditions.
@@
In the TF24 system, the pseudo-velocity components used to construct the covariance matrix $\bm{C}_{uu}$ are obtained by projecting the vectors $u_{\bm{r}_{i}}$, measured in 12 directions, onto the global x,yx,\;y, and $z$ axes. This projection is performed by multiplying each directional component by the corresponding direction vector $\bm{R}$ as,
$\bm{u}=\begin{bmatrix}u_{x}\\ u_{y}\\ u_{z}\end{bmatrix}\approx\frac{1}{4}\bm{R}^{\!\top}\widehat{\bm{u}}_{\bm{r}},$ (27)
where $\widehat{\bm{u}}_{\bm{r}}=\left[\widehat{u}_{\bm{r}_{1}}\widehat{u}_{\bm{r}_{2}}\cdots\widehat{u}_{\bm{r}_{12}}\right]^{\!\top}$.
@@ IV Results and discussion
Deviation of pseudo quantities from the C-C reference
Before presenting the main simulation results, the validity of cardioid microphone modeling and the subsequent estimation of pseudo-pressure and pseudo-velocity using the C-C method are first examined. This validation is conducted using Case 1, which serves as a baseline problem for assessing the robustness of the proposed framework.
In this study, the directional responses of practical cardioid microphones are modeled and incorporated into a TF24 configuration. The pseudo-pressure and pseudo-velocity signals are then derived from paired microphone outputs using the C-C method. The adequacy of this modeling and processing chain is evaluated by comparing the resulting pseudo-intensity estimates with those obtained under idealized conditions.
Specifically, two configurations are considered:
  1. i)
    an array composed of ideal cardioid directivities, defined as $(1-\cos\theta)/2$, and
  2. ii)
    an array using measured cardioid directivities, which include realistic deviations in amplitude, phase, and orientation.
For both configurations, the accuracy of direction estimation by pseudo-intensity is evaluated, together with the ratio between the pseudo-intensity and pseudo-energy quantities (I/E).
The latter metric is introduced as a combined indicator that reflects not only angular estimation errors but also quantitative deviations in the reconstructed intensity magnitude. As such, it provides a more comprehensive measure of the impact of microphone directivity imperfections than angular error alone.
To systematically investigate these effects, four levels of perturbations, denoted as L0 - L3, are considered for the measured cardioid data. These levels represent increasing degrees of phase, amplitude, and angular deviations, corresponding to plausible and conservative ranges of manufacturing and calibration uncertainties. For each level, the deviations are modeled as zero-mean random perturbations characterized by the standard deviations listed in Table 1, and the resulting quantities are evaluated using 200 Monte Carlo realizations. The specific definitions of L0 - L3 are summarized in Table 1.
Table 1: Assumed levels of realistic deviations in cardioid directivity used in Case 1. The values indicate standard deviations of the imposed perturbations, and each condition is evaluated using 200 Monte Carlo trials.
Level Gain [dB] Phase [deg] Axis-misalign. [deg]
L0 0.0 0 0
L1 0.5 5 1
L2 1.0 10 3
L3 2.0 20 5
Figure 8 summarizes the validation results for Case 1, showing the frequency dependence of the performance degradation caused by realistic deviations in cardioid directivity. The upper panel presents the angular error penalty, while the lower panel shows the corresponding I/E residual penalty, both defined as the difference between the measured and ideal cardioid configurations.
For each frequency band, the angular and I/E errors were first evaluated for all incident directions using Monte Carlo simulations. The 90th percentile of the error distribution was computed for each direction, and the median across all directions was then used as a representative measure. All curves, therefore, represent conservative, direction-averaged penalties relative to the ideal cardioid case.
As shown in the figure, the error-free condition L0 exhibits negligible angular deviation in the mid-frequency range, together with I/E residuals well within 1%. With increasing perturbation levels from L1 to L3, the magnitude of the deviations increases, as expected. However, even in the most severe case considered here, the absolute errors remain small across the entire frequency range, indicating that the impact of realistic cardioid imperfections is limited in practical situations.
Based on these results, the proposed modeling of practical cardioid directivities and the associated C-C-based processing are considered sufficiently robust. In the following, simulation results incorporating measured cardioid characteristics are therefore presented without further modification.
Figure 8: Numerical simulation results of frequency-dependent performance degradation caused by realistic deviations in cardioid directivity (Table 1), evaluated as the difference between modeled and ideal cardioid configurations. The upper panel shows the angular error penalty, and the lower panel shows the corresponding I/E residual penalty. Each curve represents the direction median of the Monte Carlo 90th-percentile error.
@@
Figure 9: Case 1 (numerical simulation): Diffuseness indices based on the I/E method: for Afmt and Fibo64, the results show $\Psi_{\mathrm{IE}}$; for TF24, $\Psi_{\mathrm{AVE}}$ is presented.
The Afmt results exhibited relatively large values in the low-frequency region and above 4 kHz, while the TF24 array also showed slightly elevated values of below 0.1 at a wide frequency range. In contrast, the Fibo64 array, which employs 64 microphones, yielded values close to zero across all frequency bands.
For the eigenvalue-based indices, although the absolute magnitudes differed across methods, all 2,520 incident directions yielded values of zero up to three decimal places when averaged over the nine octave bands from 63 Hz to 16 kHz; therefore, these results are not shown in the figure.
@@
Directional accuracy of the active-intensity vector was also compared. The results are shown in Fig. 10. Although diffuseness and the angular error in intensity detection are entirely different quantities, the overall trend of the results is similar to that shown in Fig. 9. For the Afmt system, the estimation accuracy is low in both the low- and high-frequency ranges, and the frequency band in which the correct incident direction can be detected is limited to approximately 500 Hz to 2 kHz. In contrast, the Fibo64 system achieves almost error-free direction detection across the entire frequency range. The TF24 array employing cardioid microphones demonstrates angular detection accuracy comparable to that of the Fibo64 configuration. Specifically, the error remains within 1 below 4 kHz, effectively matching the performance of Fibo64 in this frequency range. Although a moderate increase in error is observed at higher frequencies, it does not exceed 2, suggesting that TF24 maintains robust directional performance across a wide frequency range. Considering that these values represent the average error over 2,520 incident directions, the overall performance can be regarded as practically sufficient.
Figure 10: Case 1 (numerical simulation): Mean detection error of the incident angle estimated from acoustic intensity obtained by each microphone array: Average detection error over 2520 incident directions.
@@
Figure 11 shows the results of the COMEDIE-based diffuseness index, $\Psi_{\mathrm{COM}}$, for the TF24 system. The results closely follow the theoretical value of $1-\eta$ from the low to mid frequencies. All indices exhibit an approximately linear relationship, and the correlation between $1-\eta$ and the diffuseness estimated in each frequency band is very high. Because the indices are therefore difficult to distinguish in direct comparison, we quantified their deviation more explicitly by calculating the absolute difference between the diffuseness value and $1-\eta$, and taking the maximum of this difference for each index.
Figure 12 presents the results for (a) Afmt, (b) Fibo64, and (c) TF24. For each configuration, the intensity-based diffuseness measures $\Psi_{\mathrm{IE}}$ (or $\Psi_{\mathrm{AVE}}$ in the TF24 case), the coefficient-of-variation-based measure $\Psi_{\mathrm{CV}}$, and the eigenvalue-based measure $\Psi_{\mathrm{COM}}$ are shown.
For the Afmt configuration, the intensity-based approach exhibits large deviations at both low and high frequencies. In contrast, the error associated with the eigenvalue-based measure $\Psi_{\mathrm{COM}}$ remains small over the entire frequency range. The $\Psi_{\mathrm{CV}}$ measure also shows stable behavior, with deviations below 0.1 across all frequency bands.
The same overall tendency is observed for the Fibo64 and TF24 configurations. In all cases, $\Psi_{\mathrm{COM}}$ yields the smallest error, $\Psi_{\mathrm{CV}}$ remains consistently below 0.1, and the intensity-based measures $\Psi_{\mathrm{IE}}$ or $\Psi_{\mathrm{AVE}}$ exhibit intermediate performance.
Figure 11: Case 2 (numerical simulation): Example of the variation in diffuseness when the beam ratio $\eta$ is varied from 0 to 1 in steps of 0.05. Results are obtained for the TF24 array using the eigenvalues of the covariance matrix and the COMEDIE, i.e., $\Psi_{\mathrm{COM}}$. The numbers in the figure indicate $\eta$; in the ideal case, the diffuseness corresponds to $1-\eta$.
Figure 12: Case 2 (numerical simulation): Maximum absolute deviation between the beam ratio $1-\eta$ and the estimated diffuseness. Results are shown for (a) Afmt, (b) Fibo64, and (c) TF24 configurations. For each case, the intensity-based diffuseness measures $\Psi_{\mathrm{IE}}$ (or $\Psi_{\mathrm{AVE}}$ for TF24), the coefficient-of-variation-based measure $\Psi_{\mathrm{CV}}$, and the eigenvalue-distribution-based measure $\Psi_{\mathrm{COM}}$ are compared as a function of frequency.
@@
Figure 13: Case 3 (numerical simulation): Diffuseness for the interference field formed by two plane waves at various incident angles at $1\text{\,}\mathrm{k}\mathrm{H}\mathrm{z}$. (a) Afmt, (b) Fibo64, and (c) TF24. The results for $\Psi_{\mathrm{IE}}$, $\Psi_{\mathrm{PR}}$, and $\Psi_{\mathrm{COM}}$ are shown. For reference, in (a), the results of $\Psi_{\mathrm{PR}}$ and $\Psi_{\mathrm{COM}}$ obtained using all four FOA components ($p$ and $\bm{u}$) are also presented. In (b), the results of $\Psi_{\mathrm{COM}}$ calculated from the HOA coefficients (up to 4th) are additionally shown.
@@
Figure 14: Diffuseness calculated from numerical simulations based on measured impulse responses, obtained by mixing impulse responses measured in an anechoic room and a reverberation room with various energy ratios. The energy ratio of the anechoic component was varied from 0 to 1 in increments of 0.1. The left panels show the results for the Afmt array, while the right panels show those for the TF24 array. The upper panels correspond to the intensity-energy (I/E) method, and the lower panels show the eigenvalue-based results using the COMEDIE index.
@@ V.1 Concept
Finally, the trends obtained from the simulations were experimentally verified using measured data. Two sound fields were considered: an anechoic chamber and a reverberation chamber. As expected, diffuseness is nearly zero in the anechoic chamber and close to unity in the reverberation chamber. However, unlike the numerical simulations, neither environment can be regarded as perfectly ideal, and it is difficult to ensure that the measured results precisely match these theoretical values. Therefore, to enable a meaningful verification, we adopted an approach analogous to Case 2 of the benchmark simulations: the impulse responses measured in the anechoic and reverberation rooms were used to represent the “beam” and “diffuse” components, respectively. By mixing these two responses at arbitrary energy ratios, we examined the linearity between the prescribed ratio and the resulting diffuseness.
@@ V.2 Conditions
For the Afmt array, an AMBEO VR Mic was used, and for the TF24 array, DPA 2012 microphones were employed with a spacing of 20 mm between each microphone pair. The array was mounted on a rotatable jig (a camera tripod head) that allowed precise rotations in 45° increments around a point located 10 mm above the estimated center.
@@ V.3 Results
The results are shown in Fig. 14. The left panels correspond to the Afmt array, and the right panels to the TF24 array. The upper panels show the diffuseness obtained using the intensity-energy (I/E) method, while the lower panels present the diffuseness $\Psi_{\mathrm{COM}}$ calculated from the eigenvalues of the covariance matrix constructed only from the velocity components of the FOA signals.
In this simulation, the $\Psi_{\mathrm{AVE}}$ index obtained with the TF24 array exhibits an overall stable and approximately linear variation with respect to the mixing ratio. In contrast, the I/E-based results for the Afmt array show relatively high diffuseness values above 4kHz, while the eigenvalue-based $\Psi_{\mathrm{COM}}$ for Afmt displays a slightly biased behavior in the low-frequency range. Furthermore, for the TF24 array, the $\Psi_{\mathrm{COM}}$ index also deviates from a linear trend at 8kHz and 16 kHz under the present simulation conditions.
The fact that the diffuseness does not reach exactly unity when $\eta=0$ (reverberation room only) and does not converge to zero when $\eta=1$ (anechoic room only) is attributed to both the non-ideal characteristics of the measured sound fields and the inevitable experimental errors, so the absolute values should be interpreted with appropriate caution.
Figure 15: The maximum absolute difference between the calculated diffuseness and $1-\eta$ for each index, in which the diffuseness was calculated from numerical simulations based on measured impulse responses